OpenStudy (anonymous):

Solve for b: (2b)/5+b/10-4 =6+3b-15

OpenStudy (judygreeneyes):

Multiply by 10 to get rid of the fractions 4b + b - 40 = 60 + 30b - 150 5b - 40 = 30b - 90 subtract 5b from both sides, and add 90 to both sides 50 = 25b b = 50/25 = 2